OpenStudy (anonymous):

Need help with system: -7x-8y=9 -4x+9y=-22

OpenStudy (radar):

4(-7x-8y=9) = -28x-32y=36 7(-4x+9y=-22) = -28x+63y=-154 Subtract: 0x-95y=190 Solve for y: y=-2 Substitute -2 for y in original equation and solve for x.

OpenStudy (anonymous):

thanks